OpenStudy (anonymous):

solve by completing the square 12x^2+25x=7

OpenStudy (anonymous):

Hi this is the first time on this site, I am desperate on learning it

OpenStudy (watchmath):

\(12(x^2+\frac{25}{12}x)=7\) \(12(x^2+\frac{25}{12}x)+12\cdot (25/24)^2=7+12\cdot(25/24)^2\) \(12(x+(25/24) )^2=7+12\cdot(25/24)^2\) \((x+(25/24))^2=(7/12)+(25/24)^2\) you can continue from there :)

OpenStudy (anonymous):

12x^2 25x 7 ------ + ---- = --- 12 12 12 25x 25 7 25 x^2 + ---- + ---- = --- + ---- (12)2 12 12 24 (x+25/24)^2=13/8 x+25/24 = sqrt(13/8) x=sqrt(13/8)-25/24 you can use calculator

OpenStudy (anonymous):

hope that helps and simplify further what has watchmath had started up in there

OpenStudy (anonymous):

dlugo did you understand the solution here?

OpenStudy (anonymous):

oooops here is a better one.... 12x^2 25x 7 ------ + ---- = --- 12 12 12 25x 25 7 25 x^2 + ---- + ---- = --- + ---- 12 2*12 12 24 (x+25/24)^2=13/8 x+25/24 = sqrt(13/8) x=sqrt(13/8)-25/24 you can use calculator